# System Identification of Neural Systems: Going Beyond Images to Modelling Dynamics

### v1

@@ 4.4 Fine-grained analysis
In this section, we conduct a fine-grained analysis that goes beyond families of models. We start with studying two stream vs. single stream architectures across three video understanding datasets. Figure 3 (a) shows clearly that the two stream architectures have better ability to model the visual cortex than single stream ones in the low level regions. Additionally, they show either better or on-par scores with respect to the single stream in the high level regions. We then discuss the self-supervised learning results that showed worse regression scores in the families of models comparison with respect to full supervision. Towards this end, we investigate OmniMAE variants (i.e., self supervised and finetuned with full supervision) and TimeSformer. Figure 3 (b) shows that fully supervised models gives better scores than the self-supervised ones across the nine regions. Furthermore, we investigate which model is better at predicting the visual cortex responses. Figure 3 (c) shows that both SlowFast, a two-stream architecture, and MViT, a multiscale vision transformer, are the best in modelling the visual cortex across the nine regions. SlowFast which is a convolutional approach is better on average than MViT in the early-mid regions except the last three high-level regions (i.e., FFA, STS and PPA). These last three regions MViT shows either on-par or better regression scores. This might relate to the previous findings on how the transformers family are generally lower than the convolutional one in the early regions. However, note that MViT is equipped with multiscale processing making it better in capturing these orientations and frequencies than other transformers. Moreover, we see self-supervised models (i.e., stMAE and OmniMAE) lagging behind fully supervised ones. Additional fine-grained analysis is provided in Appendix A.3 and A.4.
Finally, we demonstrate the layer contribution of the studied video understanding models across different regions to have a better understanding of the hierarchical nature of biological neural systems. We show the layer contribution to the regression according to the regularized encoding technique in Section 3.2. Figure 4 shows the layer contribution for three video understanding models; MViT, I3D and SlowFast (Slow and Fast branches). It clearly demonstrates that across the four early layers in these networks, there is a higher contribution to the early and mid-level regions (V1-4), and the opposite occurs as we go deeper.
4.5 Summary
Here we summarize our findings: (i) our results in the simulated experiments show that system identification between image and video understanding models is attainable to a certain level. (ii) We show that video understanding models are better in regressing the visual cortex responses than image ones. (iii) We show that convolutional models predict better the early-mid regions than transformer based ones. (iv) We show that multiscale ViT (MViT) tends to perform similar to convolutional models in early-mid regions. (v) We show that two-stream video understanding models perform better than their single stream counter-part. (vi) Finally, we show that models trained with full supervision tend to surpass the ones trained in a self supervised manner.
5 Conclusion
This paper has provided a large-scale study of video understanding models with more than 30 models from a neuroscience perspective. We have shown the feasibility of system identification for the modelling scheme comparing single image vs. video understanding models. Moreover, we have shown in neural encoding experiments that modelling dynamics should be considered when comparing biological neural system responses to deep networks. Finally, we have provided insights on different families of models and fine-grained analysis considering the modelling scheme (video vs. single image), the supervision signal, and the architecture (convolutional vs. transformer based).
Appendix A Additional Results
A.1 Additional Simulated Experimental Results
Although our focus in the system identification is on the ability to differentiate image vs. video understanding models, we provide additional results for other families of models. In Figure 5 (a-c), we show the simulated results comparing convolutional vs. transformer based models for three target models MViT-B 16x4, I3D ResNet-50 and ViT-B, respectively. For the target model I3D, the figure clearly shows that convolutional models are better predictors than transformer-based ones with statistical significance across all regions in the visual cortex. The target model MViT-B shows a surprising result, that is confirming with previous findings in the real experiments as well as detailed in Section A.3, where convolutional models are better in regressing the multiscale variant than transformer-based ones. This might be explained by the fact that the multiscale ViT tends to act similarly to the convolutional models when predicting early-mid regions of the visual cortex. However, for ViT-B both are comparable and we leave it for future work to explore the reason behind this. Additionally, we report the fully vs. self supervised models for the three target models in Figure 5 (d-f). It again shows that fully supervised is capable of modelling both MViT and I3D better than the self-supervised ones (d,e). However, for the ViT-B target (f) it shows both are comparable. We believe the comparable results of the ViT-B target model might be related to having less fully supervised transformer models than the fully supervised convolutional ones since half of the transformer source models are self-supervised. For future work, we will focus on exploring the ViT-B target model by using additional fully-supervised transformer-based models.
Figure 5: Simulated experiments showing regression scores as Pearson’s correlation coefficient of: (a-c) convolutional (blue) vs. transformer (red) model families on all target models MViT-B 16x4, I3D ResNet-50 and ViT-B respectively, (d-f) fully (blue) vs. self supervised (red) model families on all target models MViT-B 16x4, I3D ResNet-50 and ViT-B respectively. Statistical significance is shown at the bottom as ‘ns’ not significant, ‘*,**,****,**,**** , * * , * * *’ significant with p-values $<0.05,0.01,0.001$, respectively.
A.2 Additional Real Experimental Results
We add a study of the model subfamilies in terms of (a) convolutional vs. transformer based models and (b) fully vs. self supervised models, but focused on video understanding models only excluding models trained with single images. As shown in Figure 6 (a) it shows consistently that convolutional based models perform better in early layers as found earlier, where the first three regions show statistically significant results. Moreover, Figure 6 (b) shows models trained in a self-supervised
Figure 6: Real experiments showing regression scores as Pearson’s correlation coefficient of model families on brain fMRI data with focus on video understanding models. (a) comparison of convolutional vs. transformer-based amongst video based models. (b) comparison of fully vs. self supervised amongst video understanding models. Statistical significance is shown in the bottom as ‘ns’ not significant, ‘*,**,****,**,**** , * * , * * *’ significant with p-values $<0.05,0.01,0.001$, resp.
learning manner tend to be worse than fully supervised ones. However, note that these results are using a small number of self-supervised learning video understanding models. Thus, we leave it for future work to expand on this further.
@@ A.2 Additional Real Experimental Results
Figure 7: Fine-grained analysis comparing instances of Convolutional (i.e., Slow) vs. Transformer (i.e., TimeSformer) based models both trained on SSV2.
A.3 Additional Fine-grained Analysis
@@
In Fig. 7 we compare instances of convolutional (i.e., Slow) and transformer (i.e., TimeSformer) based models. It clearly shows that especially across the early-mid regions in the brain (i.e., V1-4), the convolutional model tends to provide better regression scores. This confirms our previous insight that convolutional models are better at capturing orientation and frequencies because of their early local connections. In later regions in the brain (i.e., LOC, EBA, FFA, STS, PPA) convolutional model (i.e., Slow) tends to act on-par or less than the transformer-based model (i.e., TimeSformer). Moreover, we analyze the best and worst across the models trained on single images vs. videos across both the simulated and real experiments. For the simulated experiments, Table 3,4 and  5 shows the worst (Min) and best (Max) predictors from each category (Image vs. Video) for target models I3D ResNet-50, ViT-B-16 and MViT-B 16x4, respectively. For I3D and ViT, it conveys a simpler message that the best regressors are built with features extracted from architectures that exhibit higher similarity to the target model (i.e., convolutional/transformer) from both Video and Image understanding families. However, for MViT looking at Table 5 we see surprisingly the best in the Image understanding models family is ResNet-50. Although this might seem counterintuitive, yet MViT model with the addition of multiscale tends to perform similar to convolutional models not only in the simulated but also in the real experiments in the early-mid regions as shown in Fig. 2 (b), where the highest transformer based model predicting early-mid regions of the brain is MViT. Hence, we believe this might be the reason behind ResNet-50 being the best from the Image understanding models family. Additionally, we show the best and worst predictors in the real experiments with the visual cortex regions as the target in Table 6.
Table 3: Fine-grained analysis of simulated experiments results in Fig. 1 (c) with target model I3D ResNet-50 showing the worst (Min) and best (Max) source model from image & video understanding models. B1-7: Blocks in the target model.
@@
B1 stMAE SlowFast-R101 Dino ResNet-18
B2 stMAE C2D Dino ResNet-50
B3 stMAE C2D Dino ResNet-50
B4 stMAE C2D ViT-B-16 ResNet-18
B5 stMAE R(2+1)D ViT-B-16 ResNet-50
B6 stMAE C2D ViT-B-16 ResNet-18
B7 stMAE SlowFast-R101 ViT-B-16 ResNet-50
@@
B1 stMAE OmniMAE-B-Fine Dino ViT-B-32
B2 stMAE OmniMAE-B-Fine Dino ViT-B-32
B3 stMAE OmniMAE-B-Pre Dino ViT-B-32
B4 X3D-L MViT-B-32x3 Dino ViT-B-32
@@
B1 stMAE MViT-B-32x3 ViT-B-16 ResNet-50
B2 stMAE MViT-B-32x3 ViT-B-16 ResNet-18
B3 stMAE MViT-B-32x3 ViT-B-16 ResNet-50
B4 stMAE MViT-B-32x3 ViT-B-16 ResNet-50
B5 stMAE MViT-B-32x3 Dino ResNet-50
@@
LOC OmniMAE-B-Pre MViT-B-32x3 ViT-B-16 ResNet-50
EBA OmniMAE-B-Pre SlowFast-R101 ViT-B-16 ResNet-50
FFA OmniMAE-B-Pre SlowFast-R50-8x8-Char ViT-B-16 ResNet-50
STS OmniMAE-B-Pre MViT-B-32x3 ViT-B-16 ResNet-50
@@
It shows both SlowFast and MViT are the best predictors from the Video understanding family across the brain regions, with the SlowFast better at early regions similar to our previous findings as a two-stream and convolutional variant. It also shows ResNets to be the best from the Image understanding family.
A.4 Statistical Significance Results
@@
Tables 7,8 shows the pairs of models from Fig. 3 (a) [single-stream vs two-stream] and Fig. 3 (b) [fully-supervised vs self-supervised] that exhibited a statistically significant result compared to each other. Table 7 confirms that two-stream models surpass the single-stream counterpart, across 8 of the 9 visual cortex regions, with a statistically significant result. The table specifically shows that the superiority of the two-stream is independent of the training dataset. In 5 of 9 brain regions, the two-stream models were superior compared to single-stream models at two different model versions that were matched based on their training dataset (K400 and Charades). On the other hand, Table 8 confirms that fully-supervised (i.e., OmniMAE Fine and TimeSformer) surpass the self-supervised counterpart (i.e., OmniMAE Pre) across all the visual cortex regions with a statistically significant result when the three models share the same architecture base (ViT-B) and training dataset (SSV2). The table, in addition to Fig. 3 (b), shows that TimeSformer (trained solely using full-supervision) achieved the highest regression scores followed by OmniMAE Fine (trained using both self- and full-supervision) and finally OmniMAE Pre (trained solely using self-supervision). This result shows that full-supervision training is better in predicting the visual cortex responses. Table 9 shows the significance results between video understanding pairs of models. As shown in Table 9 and Fig. 3 (c), SlowFast is statistically better than I3D in 7 of the 9 brain regions, I3D is statistically better than stMAE in 5 brain regions including 4 early regions of the visual cortex (V1 to V4), while SlowFast is not statistically different than MViT in any of the regions.
A.5 Comparison to state-of-the-art methods on Algonauts Benchmark
The focus of this work is to study video understanding models from a neuroscience perspective, through a large-scale comparison of state-of-the-art deep video understanding models to the visual cortex recordings.
@@ A.5 Comparison to state-of-the-art methods on Algonauts Benchmark
Table 7: Statistical Significant of Slow vs. SlowFast (Fig. 3a), showing the pairs of models (.,.) that exhibited statistical significance. K400: Kinetics 400, Char: Charades that were used as training datasets.
@@
V2
SlowFast-K400, Slow-K400
SlowFast-Char, Slow-Char
V3
SlowFast-K400, Slow-K400
SlowFast-Char, Slow-Char
@@
LOC
SlowFast-K400, Slow-K400
SlowFast-Char, Slow-Char
EBA
SlowFast-K400, Slow-K400
SlowFast-Char, Slow-Char
FFA
SlowFast-K400, Slow-K400
SlowFast-Char, Slow-Char
@@
Table 8: Statistical Significant of Self-supervised vs. Fully Supervised (Fig. 3b), showing the pairs of models (.,.) that exhibited statistical significance.
@@
V1
OmniMAE Pre, OmniMAE Fine
OmniMAE Pre, TimeSformer
V2
OmniMAE Pre, OmniMAE Fine
OmniMAE Pre, TimeSformer
V3
OmniMAE Pre, OmniMAE Fine
OmniMAE Pre, TimeSformer
V4
OmniMAE Pre, OmniMAE Fine
OmniMAE Pre, TimeSformer
LOC
OmniMAE Pre, TimeSformer
@@
PPA
OmniMAE Pre, OmniMAE Fine
OmniMAE Pre, TimeSformer
Table 9: Statistical Significant of video understanding models (Fig. 3c), showing two pairs of models (.,.) that exhibited statistical significance.
@@
V1
SlowFast-R50-8x8, I3D
stMAE, I3D
V2
SlowFast-R50-8x8, I3D
stMAE, I3D
V3
SlowFast-R50-8x8, I3D
stMAE, I3D
V4
SlowFast-R50-8x8, I3D
stMAE, I3D
LOC
SlowFast-R50-8x8, I3D
stMAE, I3D
EBA
SlowFast-R50-8x8, I3D
PPA
SlowFast-R50-8x8, I3D
Nonetheless, we provide further comparison to the winner of the Algonauts challenge in 2021. Note that we do not have access to the test set and the challenge has been closed with the accompanying evaluation server. Therefore, we evaluate on the validation set instead of the four folds that were used throughout the above experiments and evaluate our top performing models. Figure 8 shows the winning model results on the validation set, retrieved from the original report, Yang et al. (2021) (Fig. 4 in the report), for their I3D and combined model. It is important to note that our models’ backbones were not fine-tuned during the regression, since our original question is not to win the challenge but to compare video understanding models as is, without fine-tuning. On the other hand, the winning entry used an ensemble of models with multiple modalities (i.e., audio and optical flow) and finetuned their backbones to the downstream task. Hence, it is unfair comparison with our reported models. Nonetheless, our best models perform comparable without different modalities, ensemble or finetuning of the backbones. Therefore, our experimental setup is sufficient to conclude meaningful insights about these deep video understanding models.
V1V2V3V4LOCEBAFFASTSPPA0.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Visual Cortex RegionsRegression ScoreI3DSlowFastMViTI3D-RGB (SOA)Combined (SOA)
Figure 8: Comparison to state-of-the-art model winning the Algonauts challenge 2021 but evaluated on the validation set (i.e., I3D (SOA), Combined (SOA)). Winner results retrieved from their original report, Yang et al. (2021).

### v2

@@ B.3 Additional real experimental results
Figure 12: Fine-grained analysis comparing instances of Convolutional (i.e., Slow) vs. Transformer (i.e., TimeSformer) based models with different training datasets (SSv2 and K400).
B.4 Additional fine-grained analysis
In this section, we provide additional fine-grained analysis for both real and simulated environments. First, we demonstrate the layer contribution of the studied video understanding models across different regions to have a better understanding of the hierarchical nature of biological neural systems. We show the layer contribution to the regression according to the regularized encoding technique in Section 3.3. Figure 11 shows the layer contribution for three video understanding models; MViT, I3D and SlowFast (Slow and Fast branches). It clearly demonstrates that across the four early layers in these networks, there is a higher contribution to the early and mid-level regions (V1-4), and the opposite occurs as we go deeper.
Second, in Fig. 12 we compare instances of convolutional (i.e., Slow) and transformer (i.e., TimeSformer) based models. It clearly shows that especially across the early-mid regions in the brain (i.e., V1-4), the convolutional models tend to provide better regression scores. This confirms our previous insight that convolutional models are better at capturing orientation and frequencies because of their early local connections. In later regions in the brain (i.e., LOC, EBA, FFA, STS, PPA) convolutional model (i.e., Slow) tends to act on-par or less than the transformer-based model (i.e., TimeSformer).
@@ B.4 Additional fine-grained analysis
Third, we analyze the best and worst across the models trained on single images vs. videos across both the simulated and real experiments. For the simulated experiments in Fig. 2, Table 3,4,5 and  6 show the worst (Min) and best (Max) predictors from each category (Image vs. Video) for target models MViT-B 16x4, ViT-B-16, ResNet-50, and I3D ResNet-50, respectively. For I3D, ResNet-50 and ViT, the results convey a simpler message that the best regressors are built with features extracted from architectures that exhibit higher similarity to the target model (i.e., convolutional/transformer) from both Video and Image understanding families. However, for MViT looking at Table 3 we see the best in the Image understanding models family is ResNet-50 which aligns with our previous insight that MViT tends to behave similar to convolutional models especially in the early layers.
Additionally, we show the best and worst predictors in the real experiments in Fig. 3 with the visual cortex regions as the target in Table 7. It shows both SlowFast and MViT are the best predictors from the video understanding family across the brain regions, with the SlowFast better at early regions similar to our previous findings as a two-stream and convolutional variant. It also shows ResNets to be the best from the image understanding family.
Fourth, we conduct an experiment comparing a randomly initialized network with respect to a trained model. The motivation behind this experiment is to confirm that our neural encoding is indeed capturing the learned representations in deep networks beyond random weights. Figure 13(a) shows SlowFast trained weights vs. random weights to confirm that the random weights provide worse correlation coefficient than the trained ones. This is indicative that our results are not only emergent from the architecture, but rather the learned representations that depend on architecture, dynamics encoding, training dataset and training signal (full/self supervision). Finally, we conduct experiments where we show the centered kernal alignment scores following previous works Han et al. (2023) in Fig. 13(b). It shows that SlowFast, a video understanding model, surpasses ResNet-50, a single image understanding one, even when using another metric beyond the correlation coefficient of the regressed output.
@@
B1 OmniMAE-B-Fine MViT-B-32x3 ResNet-34 ResNet-50
B2 OmniMAE-B-Fine MViT-B-32x3 ViT-B-16 ResNet-152
B3 Slow-R50-SSv2 MViT-B-32x3 ViT-B-16 ResNet-101
B4 Slow-R50-SSv2 MViT-B-32x3 ViT-B-16 ResNet-34
B5 Slow-R50-SSv2 MViT-B-32x3 ViT-B-16 ResNet-101
@@
B1 MViT-B-32x3 OmniMAE-B-Fine ResNet-101 ViT-L-16
B2 SlowFast-R50-Char OmniMAE-B-Fine ResNet-152 ViT-B-32
B3 SlowFast-R50-Char OmniMAE-B-Fine ResNet-34 ViT-L-16
B4 Slow-R50-SSv2 TimeSformer-K400 ResNet-18 ViT-B-32
Table 5: Fine-grained analysis of simulated experiments results in Fig. 2(c) with target model ResNet-50 showing the worst (Min) and best (Max) source model from image & video understanding models. B1-7: Blocks in the target model.
@@
B1 MViT-B-32x3 SlowFast-R50-SSv2 ViT-B-16 ResNet-34
B2 MViT-B-16x4 C2D ViT-B-16 ResNet-152
B3 MViT-B-16x4 C2D ViT-B-16 ResNet-18
B4 SlowFast-R50-Char OmniMAE-B-Fine ViT-B-16 ResNet-34
B5 SlowFast-R50-SSv2 TimeSformer-K400 ViT-B-16 ResNet-101
B6 Slow-R50-SSv2 TimeSformer-K400 ViT-B-16 ResNet-101
B7 Slow-R50-SSv2 TimeSformer-K400 ViT-B-16 ResNet-101
Table 6: Fine-grained analysis of simulated experiments results in Fig. 2(d) with target model I3D ResNet-50 showing the worst (Min) and best (Max) source model from image & video understanding models. B1-7: Blocks in the target model.
Min (Video) Max (Video) Min (Image) Max (Image)
B1 MViT-B-32x3 CSN ViT-B-16 ResNet-34
B2 MViT-B-16x4 SlowFast-R101 ViT-B-16 ResNet-18
B3 MViT-B-32x3 C2D ViT-B-16 ResNet-18
B4 MViT-B-32x3 C2D ViT-L-32 ResNet-18
B5 OmniMAE-B-Fine X3D-L ViT-B-16 ResNet-50
B6 SlowFast-R50-SSv2 C2D ViT-B-16 ResNet-101
B7 Slow-R50-SSv2 SlowFast-R50-K400 ViT-B-16 ResNet-101
@@
LOC OmniMAE-B-Pre MViT-B-32x3 ViT-B-16 ResNet-101
EBA OmniMAE-B-Pre SlowFast-R101 ViT-B-16 ResNet-101
FFA OmniMAE-B-Pre SlowFast-R50-8x8-Char ViT-B-16 ResNet-101
STS OmniMAE-B-Pre MViT-B-32x3 ViT-B-16 ResNet-101
@@
V1V2V3V4LOCEBAFFASTSPPA000.10.10.10.10.20.20.20.20.30.30.30.3Visual Cortex RegionsRegression ScoreSlowFastSlowFast Random
V1V2V3V4LOCEBAFFASTSPPA000.050.050.050.050.10.10.10.10.150.150.150.15Visual Cortex RegionsCKAResNet50SlowFast
Figure 13: (a) Comparison between a trained and randomized SlowFast model. (b) Centered Kernal Alignment (CKA) for ResNet50 and SlowFast models. It shows the consistency of our results using other metrics beyond the regression score.
B.5 Statistical significance results
@@
Tables 8,9(a) shows the pairs of models from Fig. 4(a) (single-stream vs. two-stream) and Fig. 4(b) (fully-supervised vs. self-supervised) that exhibited a statistically significant result compared to each other. Table 8 confirms that two-stream models surpass the single-stream counterpart, across eight of the nine visual cortex regions, with a statistically significant result. The table specifically shows that the superiority of the two-stream models is independent of the training dataset. In five of nine brain regions, the two-stream models were superior compared to single-stream models at two different model versions that were matched based on their training dataset (K400 and Charades). On the other hand, Table 9(a) confirms that fully-supervised (i.e., OmniMAE Fine and TimeSformer) surpass the self-supervised counterpart (i.e., OmniMAE Pre) across all the visual cortex regions with a statistically significant result when the three models share the same architecture base (ViT-B) and training dataset (SSv2). These results, in addition to Fig. 4(b), show that TimeSformer (trained solely using full-supervision) achieved the highest regression scores followed by OmniMAE Fine (trained using both self- and full-supervision) and finally OmniMAE Pre (trained solely using self-supervision). It demonstrates that full-supervision better reflects the visual cortex responses. Table 9(b) shows the significance results between video understanding pairs of models. As shown in Table 9(b) and Fig. 4(c), SlowFast is statistically better than I3D in seven of the nine brain regions, I3D is statistically better than stMAE in five brain regions including four early regions of the visual cortex (V1 to V4), while SlowFast is not statistically different than MViT in any of the regions.
B.6 Additional results on inter-intra region connectivity
In this section, we provide additional ablations on the connectivity module. We compare our learned connectivity to a simple random or an identity connectivity in Tables 10 and 11. In the random connectivity model, we assign random weights (using Xavier initialization) to the fully-connected layers in the connectivity model. On the other hand, in the identity connectivity model, we assign identity weights (ones) to the fully-connected layers. The connectivity model with identity weights represents a model where each region is learning from all the regions equally. The results clearly show that our learned connectivity surpasses these lower baselines of random or identity initialized connectivity.
B.7 Comparison to state-of-the-art methods on Algonauts benchmark
This work focuses on studying video understanding models from a neuroscience perspective through a large-scale comparison of state-of-the-art deep video understanding models to the visual cortex recordings. Additionally, we propose a novel encoding approach using inter-intra region connectivity on the top of pre-trained deep neural network models to predict visual cortex voxel activity. To study the effect of our novel encoding approach, we compared the accuracy achieved by our best model (MViT-B with connectivity priors) to the state-of-the-art (SOA) accuracies achieved per region  Zhou et al. (2022). In  Zhou et al. (2022), they similarly built their encoding models on pre-trained deep learning models using the layer-weighting approach and their best model varied per brain region. Accordingly, we compare our best model results with their best model per region as shown in figure 14. It shows that our novel encoding approach achieved on average higher results across all the regions except for PPA in which it is on par. This comparison shows the accuracy enhancement achieved compared to the state of the art by incorporating connectivity priors to features extracted from pre-trained deep learning models. Note that our approach has a stronger form of cross validation as we take the mean of different subjects and we compute statistics across four folds from different videos. As such, our comparison to the baseline with no connectivity that follows Zhou et al. (2022) in Fig. 5(a), better reflects our improvement gains. In future work, we will further explore the effect of integrating connectivity priors with end-to-end fine-tuning of deep learning models and ensemble approaches.
@@ B.7 Comparison to state-of-the-art methods on Algonauts benchmark
Table 8: Statistical Significant of Slow vs. SlowFast (Fig. 4(a), showing the pairs of models (.,.) that exhibited statistical significance. K400: Kinetics 400, Char: Charades that were used as training datasets.
@@
V2 SlowFast-K400, Slow-K400 SlowFast-Char, Slow-Char
V3 SlowFast-K400, Slow-K400 SlowFast-Char, Slow-Char
@@
LOC SlowFast-K400, Slow-K400 SlowFast-Char, Slow-Char
EBA SlowFast-K400, Slow-K400 SlowFast-Char, Slow-Char
FFA SlowFast-K400, Slow-K400 SlowFast-Char, Slow-Char
@@
Table 9: (a) Statistical Significant of Self-supervised vs. Fully Supervised (Fig. 4(b)), showing the pairs of models (.,.) that exhibited statistical significance. (b) Statistical Significant of video understanding models (Fig. 4(c)), showing pairs of models (.,.) that exhibited statistical significance.
@@
V1 OmniMAE Pre, OmniMAE Fine OmniMAE Pre, TimeSformer
V2 OmniMAE Pre, OmniMAE Fine OmniMAE Pre, TimeSformer
V3 OmniMAE Pre, OmniMAE Fine OmniMAE Pre, TimeSformer
V4 OmniMAE Pre, OmniMAE Fine OmniMAE Pre, TimeSformer
LOC OmniMAE Pre, TimeSformer
@@
PPA OmniMAE Pre, OmniMAE Fine OmniMAE Pre, TimeSformer
@@
V1 SlowFast-R50-8x8, I3D stMAE, I3D
V2 SlowFast-R50-8x8, I3D stMAE, I3D
V3 SlowFast-R50-8x8, I3D stMAE, I3D
V4 SlowFast-R50-8x8, I3D stMAE, I3D
LOC SlowFast-R50-8x8, I3D stMAE, I3D
EBA SlowFast-R50-8x8, I3D
PPA SlowFast-R50-8x8, I3D
Table 10: MViT learnable connectivity results compared to the base model without connectivity and other lower baselines, i.e., random and identity connectivity.
Regions Base Model Learnable Connectivity Random Connectivity Identity Connectivity
V1 0.287 0.297 -0.001 0.090
V2 0.288 0.299 0.001 0.109
V3 0.270 0.283 0.000 0.121
V4 0.263 0.272 0.000 0.108
LOC 0.293 0.316 0.000 0.163
EBA 0.349 0.369 0.000 0.223
FFA 0.281 0.306 0.002 0.164
STS 0.204 0.232 0.001 0.108
PPA 0.194 0.218 -0.002 -0.009
Table 11: SlowFast learnable connectivity results compared to the base model without connectivity and other lower baselines, i.e., random and identity connectivity.
Regions Base Model Learnable Connectivity Random Connectivity Identity Connectivity
V1 0.294 0.301 0.001 0.101
V2 0.295 0.305 -0.001 0.121
V3 0.280 0.288 -0.001 0.129
V4 0.274 0.285 -0.002 0.127
LOC 0.296 0.311 -0.001 0.156
EBA 0.354 0.360 -0.001 0.218
FFA 0.279 0.293 -0.004 0.153
STS 0.199 0.227 0.002 0.097
PPA 0.188 0.204 $0$0.019
V1V2V3V4LOCEBAFFASTSPPA0.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4Visual Cortex RegionsRegression ScoreBest Model per Region (SOA)MViT-B Connectivity
Figure 14: Comparison of MViT-B with connectivity priors to the state-of-the-art model per region as presented in  Zhou et al. (2022).
Appendix C Limitations
In this section, we discuss limitations of our work. Our current large-scale study is limited to ten subjects watching short video clips with 1,000 videos provided. We leave it for future work to explore the collection of broader datasets with longer videos up to 5-10 minutes towards the creation of foundation models for neural encoding.
Appendix D Broader impact
Conducting large-scale study on neural encoding of biological neural systems in comparison to deep neural networks, has multiple positive societal impacts as it can be used to drive insights into the understanding and advancement of deep networks. Our study also improved our understanding of the human visual cortex and its connectivity using our learned inter-intra region connectivity model, which is useful in various applications. Applications for neural encoding have various benefits including providing a way for testing and understanding brain mechanisms and computations in addition to multiple neural interfaces applications. Since the neural encoding modelling and its capabilities is still being developed and relatively in its infancy, in addition to the datasets available being limited in size, we do not perceive any negative societal impact from such developed models.
Appendix E Assets and licences
We use the Mini-Algonauts Project 2021 Cichy et al. (2021) dataset that was provided in the challenge and its development kit which is licensed under a MIT License that allows its use without restriction. The Bold Moments Dataset (BMD) including the fMRI data and stimulus metadata we used in this study is available in the OpenNeuro database under accession code DS005165 and CC0 License  Lahner et al. (2024). Most of our image and video understanding models and their trained weights are used from Pytorch Paszke et al. (2019) and Pytorch Video Fan et al. (2021a), except for few models that were retrieved from their publicly released codes and model weights.